A stone is thrown from the level ground (as in projectile motion). It is observed that the maximum height of
its flight is equal to its horizontal range. Find the launch angle θ0.
1
Expert's answer
2020-06-04T10:04:13-0400

Let, "H" is the maximum height,"R" is the horizontal range, "u" is initial velocity and "\theta" is the projectile angle.

Now, we have given "H=R" , we have to find "\theta" .

Since, we know that

"H=\frac{u^2\sin^2(\theta)}{2g}"

And

"R=\frac{u^2\sin(2\theta)}{g}"

Thus,

"H=R\implies \sin^2(\theta)=2\sin(2\theta)\\
\implies \sin^2(\theta)=4\sin(\theta)\cos(\theta)\\
\implies \tan(\theta)=4\implies\theta=\tan^{-1}(4)\approx 76^{\circ}"